Grade 12th passGeneral Physics

What is the total displacement of a freely falling body after successive rebounds from the same place of ground before it comes to stop if e is coefficient of restitution

Jhansi Rani
8 Years agoGrade 12th pass

1 Answer

Chirag Vaswani
8 Years ago
There are 2 things here the displacement and the total distance travelled.
consider the body falling from a height h
so the height to which it will rise after 1 rebound will be h/e
after 2 rebounds = h/e2
and so on
Total distance travelled will be 
d = h(1 + 1/e + 1/e2 + …. infinite terms)
this is a G.P.
d = h(1/(1-1/e))
d = he/e-1
 
and the total displacement will be h no matter how many times it rebounds.